OpenStudy (anonymous):

How do I do this? 0_0 Function M is represented by the following description. The value of y is equal to 1 less than x. Function N is represented by the following graph. Which statement correctly compares the properties of functions M and N? A. Function M has a greater rate of change and a greater y-intercept than function N. B. Function M has a greater rate of change and a smaller y-intercept than function N. C. Function N has a greater rate of change and a greater y-intercept than function M. D. Function N has a greater rate of change and a smaller y-intercept than function

OpenStudy (anonymous):

@johnweldon1993 @mathslover @mathmale @marissalovescats

OpenStudy (anonymous):

@vishweshshrimali5 @iambatman @Luigi0210 @zepdrix @tHe_FiZiCx99 @Whitemonsterbunny17 @johnweldon1993 @mathslover @mathmale @marissalovescats

OpenStudy (johnweldon1993):

Okay so M = x - 1 What does that graph look like? Where is the y-intercept? and what is the slope?

OpenStudy (anonymous):

I'm not sure, I posted the picture of the graph...

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Right...clearly it goes through the y-axis at 0 right? and the slope appears to be -1 since some points along that line are (-1,1) (-2,2) etc... So we have M = x - 1 N = -x So it appears that N has a greater y-intercept...however it appears that M has a greater rate of change So we can assume this to be B :)

OpenStudy (anonymous):

I see. :0

OpenStudy (anonymous):

3 more?

OpenStudy (anonymous):

@johnweldon1993